# Optimal Distributed Similarity Estimation for Unitary Channels

### v1

@@ Abstract
We study distributed similarity estimation for unitary channels (DSEU), the task of estimating the similarity between unitary channels implemented on different quantum devices. We completely address DSEU by showing that, for $n$-qubit unitary channels, the query complexity of DSEU is $\Theta(\sqrt{d})$, where $d=2^{n}$, for both incoherent and coherent accesses. First, we propose two estimation algorithms for DSEU with these accesses utilizing the randomized measurement toolbox. The query complexities of these algorithms are both $\mathcal{O}(\sqrt{d})$. Although incoherent access is generally weaker than coherent access, our incoherent algorithm matches this complexity by leveraging additional shared randomness between devices, highlighting the power of shared randomness in distributed quantum learning. We further establish matching lower bounds, proving that $\Theta(\sqrt{d})$ queries are both necessary and sufficient for DSEU. Finally, we compare our algorithms with independent classical shadow and show that ours have a square-root advantage. Our results provide practical and theoretically optimal tools for quantum devices benchmarking and for distributed quantum learning.
I Introduction
@@
The engineering and physical realization of quantum computers are being actively pursued across a wide range of physical platforms [1, 2, 3]. To certify their performance, numerous protocols have been developed that compare experimentally generated quantum states or channels with known theoretical targets, including direct fidelity estimation [4, 5], random benchmarking [6, 7, 8], and quantum verification [9, 10, 11, 12]. However, a central challenge is how to directly compare unknown quantum states or channels produced on different platforms, possibly at different times and locations. This task, known as cross-platform verification, becomes especially pressing as we approach regimes of quantum advantage, where classical simulation of quantum systems becomes computationally intractable. For quantum states, the core step of cross-platform verification is distributed inner product estimation. A variety of efficient algorithms have been developed for this learning task [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. Notably, for $n$-qubit quantum states, it has been shown that the sample complexity of completing this task is $\Theta(\sqrt{d})$ [15], where $d=2^{n}$.
A natural next question is how to efficiently estimate the similarity between two unitary channels. We term this learning task as distributed similarity estimation for unitary channels (DSEU). Concretely, given two $n$-qubit unitaries $U$ and $V$ implemented on separate quantum devices, the goal of DSEU is to estimate their similarity using only local quantum operations and classical communication (LOCC). This task has attracted significant attention in several recent works [24, 25, 23]. DSEU has important applications in quantum information theory, including circuit equivalence checking [26, 27, 28] and quantum channel benchmarking [29, 30, 31, 32]. Despite its importance, both lower bound and the optimal algorithm for completing DSEU remain unknown.
@@
In this work, we completely address DSEU by proving the lower bound and the corresponding optimal algorithm for both incoherent and coherent accesses. First, based on the randomized measurement toolbox [33], we provide two algorithms to accomplish DSEU with these two accesses. We show that both algorithms require only $\mathcal{O}(\sqrt{d})$ queries to the unitary channels on each device. We then establish matching lower bounds, proving that any algorithm for DSEU must make at least $\Omega(\sqrt{d})$ queries, even using multi-round LOCC, arbitrarily large ancillary systems, and adaptive state preparation and measurement (SPAM) settings. Together, these results demonstrate that our algorithms are optimal and that the query complexity of completing DSEU is $\Theta(\sqrt{d})$. We also compare our algorithms with independently performing classical shadow for unitary channels [34]. We show that independent classical shadow require $\mathcal{O}(d)$ queries to the unitary channels, whereas our protocols achieve a quadratic improvement. This square-root advantage arises from the shared randomness of SPAM settings.
II Problem Setup
Let $\mathcal{H}$ be the Hilbert space of an $n$-qubit system with dimension $d=2^{n}$. Let $\mathcal{U}(\cdot)=U(\cdot)U^{\dagger}$ and $\mathcal{V}(\cdot)=V(\cdot)V^{\dagger}$ be two unknown unitary channels implemented on different (possibly distant) quantum devices. The goal of DSEU is to estimate the similarity measure $\operatorname{Tr}^{2}[U^{\dagger}V]/d^{2}\in[0,1]$ up to additive error $\varepsilon\in(0,1)$ only using LOCC. In the following, we describe the learning models relevant to this task, focusing on two aspects: learning access and shared randomness.
@@ II Problem Setup
Figure 1: Learning models for distributed similarity estimation of unitary channels. (a) Learning access: (i) Incoherent access: the unknown channel is queried once before each measurement. (ii) Coherent access: the unknown channel is queried $T$ times before each measurement, with arbitrary intermediate quantum channels $\{\mathcal{C}_{t}\}$. (b) Shared randomness: (i) Devices share randomness used in both state preparation and measurement (SPAM) settings. (ii) Devices share randomness only for state preparation. (iii) Devices have no shared randomness.
@@
Following the classifications in [35, 36, 37, 38, 34, 39], we consider two types of learning access: incoherent and coherent accesses, as shown in Fig. 1(a). For clarity, we describe the access model for one device. (i) Incoherent access, also known as learning without quantum memory or without control, allows the learner to use the unknown channel only once before measurement. That is, in the $t$-th query, the learner prepares an arbitrary quantum state $\rho_{t}$, applies either $(\mathcal{I}\otimes\mathcal{U})(\rho_{t})$ or $(\mathcal{I}\otimes\mathcal{V})(\rho_{t})$, and then performs an arbitrary quantum measurement on the output state, which is described by a positive operator-valued measure (POVM). Here $\mathcal{I}$ is the identity channel acting on an ancillary system. (ii) Coherent access, also known as learning with quantum memory or with control, permits multiple queries of the unknown channel before measurement. In each experimental round, the learner may interleave uses of the unknown channel with arbitrary quantum channels. Therefore, the resulting output state before the measurement has the form
$(\mathcal{U}\otimes\mathcal{I})\circ\mathcal{C}_{T-1}\circ(\mathcal{U}\otimes\mathcal{I})\circ\cdots\mathcal{C}_{1}\circ(\mathcal{U}\otimes\mathcal{I})(\rho),$ (1)
with an analogous expression for $\mathcal{V}$, where $\rho$ is the initial state and $\{\mathcal{C}_{i}\}$ are arbitrary quantum channels. An arbitrary POVM is then applied to the final state.
@@
In addition to learning accesses, shared randomness plays a crucial role in randomized measurement-based distributed learning. As in many learning protocols, SPAM settings are essential for learning channels. In the framework of the randomized measurement toolbox, both initial states and measurement settings are sampled randomly to enhance practicality and sample efficiency. Thus, allowing devices to share part or all of randomness can significantly improve performance, as demonstrated in distributed inner product estimation [15]. In this work, we consider three levels of shared randomness, illustrated in Fig. 1(b): (i) shared randomness in both SPAM settings, (ii) shared randomness only in state preparation, and (iii) no shared randomness between devices. We will show that shared randomness provides clear advantages in the efficiency of distributed learning algorithms.
III Upper bound
Here, we provide two algorithms to complete DSEU and provide the corresponding query complexities.
III.1 Incoherent Access
We first consider learning with incoherent access and share the randomness of SPAM settings. Let $\{|{a}\rangle\!\langle{a}|\}_{a}$ be the computational basis. To complete DSEU, we run the following procedure $T$ times independently:
@@ III.1 Incoherent Access
    Randomly generate a pure state $|{\psi}\rangle\in\mathcal{H}$ from a state $4$-design ensemble on each quantum devices.
@@
    Apply the unitary channels $\mathcal{U}$ and $\mathcal{V}$ on each $|{\psi}\rangle$.
@@
    Measure two rotated states $\mathcal{U}(|{\psi}\rangle\!\langle{\psi}|)$ and $\mathcal{V}(|{\psi}\rangle\!\langle{\psi}|)$ in the basis $\{Q^{\dagger}|{a}\rangle\!\langle{a}|Q\}_{a}$ for $m$ times, where $Q$ is randomly sampled from a unitary $4$-design ensemble.
@@
    Recode the measurement results as $\{a_{i}\}_{i=1}^{m}$ and $\{b_{i}\}_{i=1}^{m}$, respectively. Compute
@@
    Obtain an unbiased estimator for $\operatorname{Tr}^{2}[U^{\dagger}V]/d^{2}$:
@@
This incoherent algorithm is summarized in Algorithm 1. As we can see, our incoherent algorithm is ancilla-free and non-adaptive, making it immediately implementable on near-term devices. Additionally, the corresponding query complexity is shown in the following theorem.
@@
The expectation and variance of estimator $\tilde{\omega}$ defined in Eq. (3) are given by
$\mathbb{E}\tilde{\omega}=\frac{1}{d^{2}}\operatorname{Tr}^{2}[U^{\dagger}V],\quad\mathbb{V}(\tilde{\omega})=\mathcal{O}\left(\frac{d}{m^{2}}+\frac{1}{m}+\frac{1}{d}\right).$ (4)

### v2

# Optimal Distributed Similarity Estimation of Quantum Channels

@@ Abstract
We study distributed similarity estimation of quantum channels (DSEC), a primitive for cross-platform verification where two remote quantum devices are compared by estimating the inner product of their Choi states. We show that the optimal channel query complexity of DSEC for two $d$-dimensional quantum channels is $\Theta(\max\{\sqrt{d}/\varepsilon,1/\varepsilon^{2}\})$, where $\varepsilon$ is the additive error. We first prove an information-theoretic lower bound with this scaling, which holds even in the strongest setting, allowing adaptive strategies, multiple rounds of classical communication, and coherent access with arbitrary ancillas. We then give a matching upper bound in the weakest setting, namely non-adaptive and ancilla-free incoherent access, via a randomized measurement protocol achieving this bound. Finally, we show that our protocol achieves a quadratic improvement over classical shadow baselines. Our results provide theoretically optimal and practical methods for cross-platform verification, quantum device benchmarking, and distributed quantum learning.
@@
Introduction.— The engineering and physical realization of quantum computers are being actively pursued across a wide range of physical platforms [1, 2, 3]. To certify their performance, numerous protocols have been developed that compare experimentally generated quantum states or channels with known theoretical targets, including direct fidelity estimation [4, 5], randomized benchmarking [6, 7, 8], and quantum verification [9, 10, 11, 12]. However, these protocols rely on knowledge of the theoretical target, which may not always be available. A central challenge is how to directly compare unknown quantum states or channels produced on different platforms, possibly at different times and locations. This task, known as cross-platform verification, becomes especially pressing as we approach regimes of quantum advantage, where classical simulation of quantum systems becomes computationally intractable. For quantum states, the core step of cross-platform verification is distributed inner product estimation (DIPE). A variety of efficient algorithms have been developed for this task [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. Notably, for $d$-dimensional quantum states, it has been shown that the sample complexity of completing this task is $\Theta(\max\{\sqrt{d}/\varepsilon,1/\varepsilon^{2}\})$ [15], where $\varepsilon$ is the additive error.
Building upon the success of DIPE, a natural next question is cross-platform verification of quantum channels. The corresponding core task is estimating the inner product of their Choi states using only local operations and classical communication (LOCC), which we term distributed similarity estimation of quantum channels (DSEC). DSEC has attracted significant attention in several recent works [24, 25, 23], due to important applications in quantum information theory, including circuit equivalence checking [26, 27, 28] and quantum channel benchmarking [29, 30, 31, 32]. Despite its importance, fundamental questions about DSEC remain open: What is the lower bound for completing DSEC? Can we design optimal algorithms that achieve this bound? Several existing techniques can potentially address this task, including process tomography, extension of DIPE, and classical shadow for channels. However, all of these approaches face significant limitations. Process tomography is highly resource-intensive, requiring $\mathcal{O}(d^{4})$ queries [33, 34]. Directly extending DIPE by preparing Choi states and then applying DIPE algorithms leads to $\mathcal{O}(d)$ query complexity [15] and is experimentally challenging due to the need for preparing maximally entangled states. Classical shadow for channels, while sample-efficient in many cases [35, 36, 37], is far from optimal as we will discuss later.
@@
In this work, we settle the optimal query complexity of DSEC by proving the lower bound and constructing an estimation protocol with matching upper bound based on randomized measurements [38]. First, we establish an information-theoretic lower bound, proving that any algorithm for DSEC must require at least $\Omega(\max\{\sqrt{d}/\varepsilon,1/\varepsilon^{2}\})$ queries. Remarkably, this bound holds even in the strongest setting: multi-round LOCC, arbitrarily large ancillary systems, coherent access, and adaptive strategies. We further show that this lower bound is achievable. For unitary channels, we provide a coherent algorithm with query complexity $\mathcal{O}(\max\{\sqrt{d}/\varepsilon,1/\varepsilon^{2}\})$. For general channels, we provide an algorithm requiring only non-adaptive strategies, incoherent access, and ancilla-free protocols. Surprisingly, by leveraging shared randomness between devices, this incoherent algorithm also achieves the same query complexity of $\mathcal{O}(\max\{\sqrt{d}/\varepsilon,1/\varepsilon^{2}\})$. This highlights the power of shared randomness in distributed quantum learning and shows that the query complexity of DSEC is exactly $\Theta(\max\{\sqrt{d}/\varepsilon,1/\varepsilon^{2}\})$. Finally, we compare our algorithms with independent classical shadow for channels [35, 36, 37]. We show that independent classical shadow requires $\mathcal{O}(\max\{d/\varepsilon,1/\varepsilon^{2}\})$ queries, whereas our protocols achieve a quadratic improvement. This square-root advantage arises from the shared randomness.
@@
Figure 1: Learning models for distributed channel learning. (a) Learning access: (i) Incoherent access: the unknown channel is queried once before each measurement. (ii) Coherent access: the unknown channel is queried $T$ times before each measurement, with arbitrary intermediate quantum channels $\{\mathcal{C}_{t}\}$. (b) Shared randomness: (i) for both state preparation and measurement (SPAM) settings. (ii) only for state preparation. (iii) no shared randomness.
Problem Setup.— We now formally define the DSEC problem. Let $\mathcal{H}$ be the Hilbert space of a $d$-dimensional system. Let $\mathcal{E}$ and $\mathcal{F}$ be two unknown channels implemented on different (possibly distant) quantum devices. Define the Choi operator of $\mathcal{E}$ as $J_{\mathcal{E}}\mathrel{\mathop{\ordinarycolon}}=(\mathcal{E}\otimes\mathcal{I})(|{\Phi}\rangle\!\langle{\Phi}|)$, where $\mathcal{I}$ is the identity channel and $|{\Phi}\rangle\mathrel{\mathop{\ordinarycolon}}=\sum|{ii}\rangle$ is the unnormalized maximally entangled state in $\mathcal{H}^{\otimes 2}$ [39], and similarly for $J_{\mathcal{F}}$. The goal of DSEC is to estimate the inner product $\operatorname{Tr}[J_{\mathcal{E}}J_{\mathcal{F}}]/d^{2}\in[0,1]$ up to additive error $\varepsilon\in(0,1)$ with success probability at least $2/3$. This inner product can quantify the similarity between $\mathcal{E}$ and $\mathcal{F}$ via the max fidelity [24, 40]:
$\frac{\operatorname{Tr}[J_{\mathcal{E}}J_{\mathcal{F}}]}{\max\{\operatorname{Tr}[J_{\mathcal{E}}^{2}],\operatorname{Tr}[J_{\mathcal{F}}^{2}]\}},$ (1)
where $\operatorname{Tr}[J_{\mathcal{E}}^{2}]$, $\operatorname{Tr}[J_{\mathcal{F}}^{2}]$ describe the unitarity of $\mathcal{E}$, $\mathcal{F}$. Unitarity estimation has been considered in [41], and our algorithm can also serve as a new tool for it. In the following, we describe the learning models relevant to DSEC, focusing on two aspects: learning access and shared randomness.
@@
Following the classifications in [42, 43, 44, 41, 37, 45], we consider two types of learning access, illustrated in Fig. 1(a). (i) Incoherent access, also known as learning without quantum memory: In each query, the learner prepares an arbitrary quantum state, applies the unknown channel, and immediately measures the output state. The key restriction is that the unknown channel is queried exactly once before each measurement. (ii) Coherent access, also known as learning with quantum memory: In each experimental round, the learner prepares an arbitrary quantum state, may interleave multiple queries of the unknown channel with arbitrary intermediate quantum channels, and then applies an arbitrary measurement on the final output state. This allows quantum information to be preserved across multiple channel queries.
@@
Beyond learning access, shared randomness plays a crucial role in distributed learning with randomized measurements [38]. In many learning protocols, state preparation and measurement (SPAM) settings are essential components. Within the framework of randomized measurements, both initial states and measurement settings are sampled randomly to enhance practicality and sample efficiency. In the distributed scenario, allowing devices to share part or all of this randomness can significantly improve performance, as demonstrated in DIPE [15]. We consider three levels of shared randomness, illustrated in Fig. 1(b): (i) shared randomness for both SPAM settings, (ii) shared randomness only for state preparation, and (iii) no shared randomness between devices. A key finding of this work is that shared randomness provides substantial advantages in completing DSEC.
Lower Bound.— We now establish the lower bound for completing DSEC. Our strategy is to reduce DSEC to a simpler hypothesis testing problem and analyze its query complexity. Specifically, we consider the following distinguishing problem, which serves as a restricted version of DSEC.
Problem 1.
Distinguish between the following two cases:
Two devices perform the same unitary $U$, which is a Haar random unitary;
Two devices independently perform two unitaries $U$ and $V$, which are independent Haar random unitaries.
Let $J_{U}$ and $J_{V}$ denote the Choi operators of $U$ and $V$, respectively. If a learning algorithm can estimate the value of $\operatorname{Tr}[J_{U}J_{V}]/d^{2}$, then we can use this algorithm to complete Problem 1. Hence, a lower bound for completing Problem 1 directly implies a lower bound for completing DSEC. Moreover, this lower bound also applies to estimating other commonly used distance measures between quantum channels. In the following, we analyze the query complexity required to complete Problem 1 with incoherent and coherent accesses.
To analyze the query complexity of Problem 1, we employ Le Cam’s two-point method [46], which upper bounds the success probability of hypothesis testing in terms of the total variation distance (TVD) between the probability distributions of measurement outcomes under the two hypotheses. Let $p_{1}(\ell)$ and $p_{2}(\ell)$ be the probabilities of obtaining measurement result $\ell$ after querying the channels in cases (i) and (ii), respectively. Their TVD is defined as
$\left\lVert p_{1}-p_{2}\right\rVert_{\rm TV}\mathrel{\mathop{\ordinarycolon}}=\frac{1}{2}\sum_{\ell}\left|p_{1}(\ell)-p_{2}(\ell)\right|.$ (2)
For reliable hypothesis testing, this TVD must be at least constantly large. Now, we are ready to prove the lower bounds for both learning access models.
Incoherent Access.
We first analyze the case of incoherent access, following the learning tree framework [44, 41, 47, 48, 49]. We show that incoherent access requires exponentially many queries in $\sqrt{d}$, even with multi-round LOCC, arbitrarily large ancillary systems, and adaptive SPAM settings.
Theorem 1.
Any algorithm with incoherent access, possibly utilizing multi-round LOCC, ancillary systems, and adaptive strategies, requires $\Omega(\sqrt{d})$ queries to complete Problem 1.
This theorem implies that any algorithm with incoherent access requires at least $\Omega(\sqrt{d})$ queries to complete DSEC. Notably, this scaling matches that of unitarity estimation [41], consistent with the fact that unitarity estimation is a special case of DSEC.
Proof sketch of Theorem 1.
Inspired by [41], we introduce an intermediate distribution $p_{\mathcal{D}}(\ell)$, corresponding to the probability of obtaining measurement result $\ell$ when both channels are replaced by the completely depolarizing channel $\mathcal{D}(A)\mathrel{\mathop{\ordinarycolon}}=\operatorname{Tr}[A]I/d$. By the triangle inequality, the target TVD is upper bounded by two terms: $\left\lVert p_{1}-p_{\mathcal{D}}\right\rVert_{\rm TV}$ and $\left\lVert p_{\mathcal{D}}-p_{2}\right\rVert_{\rm TV}$. We prove that each term is bounded by $\mathcal{O}(T^{2}/d)$, where $T$ is the number of queries per device. Therefore, completing Problem 1 with incoherent access requires $T=\Omega(\sqrt{d})$ queries. The detailed proof is provided in Appendix B.1. ∎
Coherent Access.
We now turn to coherent access, where the learner may interleave multiple queries with arbitrary quantum channels before measurement. In various quantum learning tasks, coherent access often provides substantial advantages over incoherent access [44, 41, 37, 45]. A central question is whether such advantages exist for DSEC. Remarkably, we prove that even with this significantly more powerful learning access, completing Problem 1 still requires $\Omega(\sqrt{d})$ queries. This result holds even with multi-round LOCC, arbitrarily large ancillary systems, and adaptive SPAM settings, highlighting the fundamental hardness of DSEC.
Theorem 2.
Any algorithm with coherent access, possibly utilizing multi-round LOCC, ancillary systems, and adaptive strategies, requires $\Omega(\sqrt{d})$ queries to complete Problem 1.
Likewise, this theorem implies that any algorithm with coherent access requires at least $\Omega(\sqrt{d})$ queries to complete DSEC. Importantly, we will show that coherent access offers no asymptotic advantage for completing DSEC: our incoherent algorithm matches this lower bound.
Proof sketch of Theorem 2.
Suppose we query $T$ times on each device. Analyzing coherent access is generally subtle due to the arbitrary quantum channels that may appear between queries. To overcome this difficulty, inspired by [50], we reformulate the problem in terms of Choi operators. Specifically, we work with the Choi operator of $U^{\otimes T}$, denoted $J_{U}^{(T)}$. This formulation absorbs all interleaving channels into an enlarged effective input system. The two probabilities $p_{1}$ and $p_{2}$ can then be expressed via the following averaged Choi operators:
$\bar{J}_{1}^{(T)}\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}_{U}\;J_{U}^{(T)}\otimes J_{U}^{(T)},\quad\bar{J}_{2}^{(T)}\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}_{U,V}\;J_{U}^{(T)}\otimes J_{V}^{(T)}.$ (3)
Subsequently, as in the proof of lower bound for incoherent access, we introduce an intermediate probability distribution $p_{a}(\ell)$, defined as the probability of obtaining measurement result $\ell$ when both channels are replaced by the approximate $T$-design channel with Choi operator $J_{a}^{(T)}$ (see definition in [50], also Appendix Eq. (29)). We again analyze the TVDs: $\left\lVert p_{1}-p_{a}\right\rVert_{\rm TV}$ and $\left\lVert p_{a}-p_{2}\right\rVert_{\rm TV}$ separately. For the first term, we directly bound it using the relative error of the approximate $T$-design channel, and the corresponding result is $\mathcal{O}(T^{2}/d)$. For the second term, we relate it to the distance between the approximate $T$-design channel and a “measure-and-prepare” map. Using tools from quantum cloning theory [51], we show that this term is also bounded by $\mathcal{O}(T^{2}/d)$. Therefore, completing DSEC with coherent access requires $T=\Omega(\sqrt{d})$ queries. The detailed proof is provided in Appendix B.2. ∎
Lower Bound for DSEC.
Combining the above results with an additional analysis of the $\varepsilon$-dependence, we establish the complete lower bound for completing DSEC. The result, proven in Appendix B.3, is as follows.
Theorem 3.
Any algorithm, possibly utilizing multi-round LOCC, ancillary systems, and adaptive strategies, requires $\Omega(\max\{\sqrt{d}/\varepsilon,1/\varepsilon^{2}\})$ queries to complete DSEC.
This theorem provides an information-theoretic lower bound that holds regardless of the learning access model.
Upper bound.— Having established the lower bound, we now present algorithms that achieve the optimal query complexity. We provide two algorithms: one for coherent access (applicable to unitary channels) and one for incoherent access (applicable to general channels).
Coherent Access.
We first consider the case where both unknown channels are unitary, i.e., $\mathcal{E}(\cdot)=U(\cdot)U^{\dagger}$ and $\mathcal{F}(\cdot)=V(\cdot)V^{\dagger}$. For this setting, we propose an algorithm based on symmetric collective measurements [52, 37]. Given a state $(s+2)$-design $\{|{\phi_{j}}\rangle\}_{j=1}^{L}$, the symmetric collective measurement $\mathcal{M}_{s}$ on $\mathcal{H}^{\otimes s}$ is defined as the following POVM:
$\mathcal{M}_{s}\mathrel{\mathop{\ordinarycolon}}=\left\{\frac{\kappa_{s}}{L}|{\phi_{j}}\rangle\!\langle{\phi_{j}}|^{\otimes T}\right\}_{j=1}^{L}\cup\left\{I-\Pi_{\rm sym}^{(d,s)}\right\},$ (4)
where $\kappa_{s}\mathrel{\mathop{\ordinarycolon}}=\binom{d+s-1}{s}$ and $\Pi_{\rm sym}^{(d,s)}$ is the projector onto the symmetric subspace of $\mathcal{H}^{\otimes s}$. Our algorithm proceeds as follows:
@@
    Randomly generate pure state $|{\psi}\rangle^{\otimes T}$ on each quantum device, where $|{\psi}\rangle$ from a state $4$-design ensemble.
@@
    Apply the unitary operators $U^{\otimes T}$ and $V^{\otimes T}$ on $|{\psi}\rangle^{\otimes T}$.
@@
    Measure $(U|{\psi}\rangle)^{\otimes T}$ and $(V|{\psi}\rangle)^{\otimes T}$ with the POVM $\mathcal{M}_{T}$. Record the measurement results on two devices as $|{\phi_{A}}\rangle$ and $|{\phi_{B}}\rangle$, and compute $\tilde{f}=|\langle\phi_{A}|\phi_{B}\rangle|^{2}$.
@@
    Return the unbiased estimator:
This algorithm requires shared randomness only for state preparation and is described in pseudocode in Algorithm 1. We summarize the query complexity of this algorithm in the following theorem, which is proven in Appendix C.1.
Theorem 4.
For unitary channels $\mathcal{E}$ and $\mathcal{F}$, the expectation and variance of estimator $\tilde{\chi}$ defined in Eq. (5) are given by
$\mathbb{E}\tilde{\chi}=\frac{1}{d^{2}}\operatorname{Tr}[J_{\mathcal{E}}J_{\mathcal{F}}],\quad\mathbb{V}(\tilde{\chi})\leq\mathcal{O}\left(\frac{d}{m^{2}}+\frac{1}{m}+\frac{1}{d}\right).$ (6)
Combined with Theorem 3, this coherent algorithm is optimal in both $d$ and $\varepsilon$ for unitary channels. However, this algorithm fundamentally requires unitary channels since the proof relies on pure output states (see Appendix C.1 for details). While random purification channels [53, 54, 34] could potentially extend this result to general channels, we pursue a different approach: we show that incoherent access already suffices to achieve optimality.
Incoherent Access.
We now present our main algorithm, which works for general channels using only incoherent access. Let $\{|{a}\rangle\!\langle{a}|\}_{a}$ be the computational basis. The algorithm repeats the following procedure $T$ times independently:
  1. 1.
    Randomly generate a pure state $|{\psi}\rangle$ from a state $4$-design ensemble on each quantum device.
  2. 2.
    Apply the unknown channels $\mathcal{E}$ and $\mathcal{F}$ on each $|{\psi}\rangle$.
  3. 3.
    Measure two output states $\mathcal{E}(|{\psi}\rangle\!\langle{\psi}|)$ and $\mathcal{F}(|{\psi}\rangle\!\langle{\psi}|)$ in the basis $\{Q^{\dagger}|{a}\rangle\!\langle{a}|Q\}_{a}$ $m$ times, where $Q$ is randomly sampled from a unitary $4$-design ensemble.
  4. 4.
    Record the measurement results as $\{a_{i}\}_{i=1}^{m}$ and $\{b_{i}\}_{i=1}^{m}$, respectively. Compute
@@
    Return the unbiased estimator:
@@
This incoherent algorithm is summarized in Algorithm 2. Additionally, the corresponding query complexity is shown in the following theorem, which is provided in Appendix C.2.
@@
For general channels $\mathcal{E}$ and $\mathcal{F}$, the expectation and variance of estimator $\tilde{\omega}$ defined in Eq. (8) are given by
$\mathbb{E}\tilde{\omega}=\frac{1}{d^{2}}\operatorname{Tr}[J_{\mathcal{E}}J_{\mathcal{F}}],\quad\mathbb{V}(\tilde{\omega})\leq\mathcal{O}\left(\frac{d}{m^{2}}+\frac{1}{m}+\frac{1}{d}\right).$ (9)